OpenStudy (anonymous):

Find a pair of factors for each number by using the difference of two squares. a. 45 b. 77 c. 112

OpenStudy (anonymous):

@snowman97,

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

can u help me in this

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

u half to pick one

OpenStudy (anonymous):

b

OpenStudy (anonymous):

ok u got it ur right b 77

OpenStudy (unklerhaukus):

45=49-4

OpenStudy (anonymous):

u half to find two squres for all

OpenStudy (unklerhaukus):

do you know what number squares to give 49?

OpenStudy (anonymous):

4

OpenStudy (anonymous):

this is my answer but my teacher mark it wrong

OpenStudy (anonymous):

45plus49-4equal7+2.7-2equal 9 times 5 77equal 81-4equal 9+2.9-2equal 11times 7 112equal 121-9equal 11plus311-3equal 14 times 8 Comment: Don't use words. Use the symbols.

OpenStudy (unklerhaukus):

\[45 = 49-4=7^2-2^2=(7-2)(7+2)=5\times9\]

OpenStudy (anonymous):

beside the word is the answer right

OpenStudy (anonymous):

wont it be 8 times

OpenStudy (unklerhaukus):

@snowman97 \(5\times8=40\)

OpenStudy (anonymous):

what about B and c @UnkleRhaukus

OpenStudy (anonymous):

wate i did that wrong

OpenStudy (unklerhaukus):

\[77=a^2-b^2=(a-b)(a+b)=\]

OpenStudy (anonymous):

40 right

OpenStudy (unklerhaukus):

can you find \(a\) and \(b\)

OpenStudy (anonymous):

my computer acting weird give me min for the answer

OpenStudy (anonymous):

77= 81 - 4= 9+ 2. 9-= 11 x 7

OpenStudy (anonymous):

@UnkleRhaukus

OpenStudy (unklerhaukus):

do you mean 77= 81 - 4= 9^2- 2^2 = (9+2)(9-2)= ?

OpenStudy (anonymous):

18

OpenStudy (unklerhaukus):

???//?

OpenStudy (anonymous):

i don't know

OpenStudy (anonymous):

he didnt give u da answer he dosnet know it

OpenStudy (anonymous):

dang dis is hard

OpenStudy (anonymous):

well do u know it

OpenStudy (anonymous):

i know it's hard

OpenStudy (anonymous):

77=81-4= 9^2- 2^2 = (9+2)(9-2)= 112

OpenStudy (anonymous):

c. 112